OpenStudy (anonymous):

Find dy/dx with implicit differentiation: e^(x/y)=2x-y

OpenStudy (anonymous):

d/dx [e^(x/y)] = d/dx [2x - y] [ d/dx [(x/y)] ] * e^(x/y) = 2 - dy/dx [1/y - (x/y^2) * dy/dx] * e^(x/y) = 2 -dy/dx now isolate dy/dx ..

OpenStudy (australopithecus):

oh I see I didnt treat x/y as a function

OpenStudy (australopithecus):

or rather y as a function in x/y

OpenStudy (anonymous):

Here we need to use the chain rule to solve the that...

OpenStudy (anonymous):

just want to check myself as well isolating dy/dx from my expression gives : [1/y - (x/y^2) * dy/dx] * e^(x/y) = 2 -dy/dx dy/dx ( 1 - (x/y^2)e^(x/y) )= 2 - [e^(x/y)]/ y dy/dx = [ 2y- e^(x/y) ]/[ (y - x/y)e^(x/y) ]

OpenStudy (anonymous):

yes same as pavanneella (i can multiply num and denom by y)

OpenStudy (anonymous):

Thanks to all!